# Winterthurerstrasse 190, 8057 Zürich, Switzerland

### 25 Mar 2026 v1

@@ Abstract
Context. The interiors of giant planets are commonly assumed to be convective and therefore adiabatic, making the adiabatic temperature gradient a key ingredient in interior and evolution models. While there are multiple numerically distinct ways to compute this gradient, their impact on inferred planetary structure and radius has not been systematically assessed.
Aims. We investigate how the numerical treatment of adiabatic temperature profiles affects inferred planetary radii and internal structure, quantifying the impact of different methods for calculating the adiabatic gradient and different forms of the temperature differential equation on static interior models.
Methods. We computed static interior models of a one Jupiter mass planet using a state-of-the-art hydrogen-helium equation of state, comparing five methods for evaluating the adiabatic gradient against a ground-truth isentropic baseline, for both the logarithmic and non-logarithmic forms of the temperature equation.
Results. The choice of numerical method significantly impacts the inferred interior structure and the radius. With the logarithmic temperature equation, central temperatures deviate by several thousand Kelvin and surface radii differ by up to 3.4%, exceeding the $\sim$ 1% precision of current giant exoplanet radius measurements threefold. The non-logarithmic form reduces deviations to below $\sim$ 1% for most methods.
Conclusions. We therefore recommend spline derivatives to evaluate the adiabatic gradient via the triple-product rule, combined with the non-logarithmic temperature equation. Finite differencing and direct use of tabulated gradients or derivatives should be avoided.
Key Words.:
planets and satellites: general, gaseous planets, interiors, method: numerical
[…]
@@ 1 Introduction
Giant planets are key objects for understanding planetary system formation and evolution, and their interior structure provides crucial constraints on their composition, thermal state, and history (Guillot 1999; Burrows et al. 2001; Fortney & Nettelmann 2010; Thorngren & Fortney 2019; Müller & Helled 2023; Helled & Howard 2026). Accurately characterising giant planet interiors therefore requires reliable structural models. A common and well-established assumption is that giant planet interiors are largely convective, leading to adiabatic (isentropic) temperature profiles (e.g., Stevenson 1982; Guillot 2005; Helled et al. 2014). The adiabatic temperature gradient thus plays a central role in both static interior models and long-term evolution calculations. However, despite being mathematically equivalent in the continuous limit, different numerical implementations of the adiabatic gradient can yield meaningfully different results, and the sensitivity of interior models to these numerical choices has not been systematically explored. In this work, we investigate and quantify the differences in the planetary interiors and radii that arise when using different methods of modelling adiabatic interiors of giant planets.
@@ 2 Methods
Giant planets are commonly modelled as a spherically symmetric object in hydrostatic equilibrium. Under these assumptions the structure equations are (e.g.., Kippenhahn et al. 2012):
@@
where $r$, $m$, $\rho$, $p$ and $T$ are the radius, mass, density, pressure and temperature, G is the gravitational constant, and $\nabla_{\rm{rad,ad}}$ are the radiative and adiabatic temperature gradients. These equations are solved together with an appropriate equation of state and boundary conditions. Equation 3 further assumes the simple case [^1] While more sophisticated models use the mixing length theory (Böhm-Vitense 1958) to determine the temperature gradient in convective regions, the adiabatic gradient still plays a key role (e.g., Kippenhahn et al. 2012). that the temperature gradient inside the planet is the minimum of the radiative or the adiabatic gradient, which is equivalent to the Schwarzschild criterion (Schwarzschild 1906). In this work, we focus solely on fully convective interiors, such that the temperature gradient is always the adiabatic one. The form of $\nabla_{\rm{ad}}$ commonly used in stellar and planetary structure calculations is the logarithmic temperature derivative with respect to pressure at constant entropy $S$ (e.g., Lamers & Levesque 2017):
@@
where $S$ is the entropy. Since the adiabatic gradient is a material property, calculating it requires an appropriate equation of state. The equation of state is generally given in a tabulated form using either the thermodynamic basis $(p,T)$ or $(\rho,T)$, and is commonly evaluated with cubic splines since derivatives can be required (for example, when using the Henyey method to solve the evolution equations (Henyey et al. 1965)). For simplicity and without loss of generality, we here focus on hydrogen-helium mixtures and do not include heavy elements.
The state-of-the-art hydrogen-helium equation of state for modelling giant planets and brown dwarfs are the tables from Chabrier & Debras (2021) (hereafter CD21). These tables provide the entropy, logarithmic entropy derivatives with respect to pressure and temperature, and the adiabatic gradient as a function of the pressure-temperature or density-temperature thermodynamic bases. There are tables for pure hydrogen and helium, or hydrogen-helium mixtures for a few values of the helium mass fraction. The hydrogen-helium mixture tables offer (among others [^2] Using Maxwell relations and the triple-product-rule, there are many ways deriving expressions for the adiabatic temperature gradient; see, e.g., Hansen & Kawaler (1994) or Equation 12 in Chabrier et al. (2019). These can be useful if the thermodynamic basis is not $(p,T)$. two relatively simple ways modelling adiabatic interiors: using the tabulated values for $\nabla_{\rm{ad}}$, or calculating it using the triple-product rule
@@ 3 Results
In this section we performed several numerical experiments with Python to determine the influence the previously mentioned numerical details have on the interiors of giant planets. We assumed a proto-solar hydrogen-helium mixture and evaluated the equations of state tables using cubic splines with scipy.interpolate.RegularGridInterpolator. The temperature equation was solved with scipy.integrate.solve_ivp using the explicit fifth-order Runge-Kutta method RK45 (additional solvers were tested in Appendix A), and the structure equations with a relaxation method (Thorngren et al. 2016; Mankovich et al. 2019; Mankovich & Fortney 2020). We assumed a planetary mass of 1 MJ and the surface boundary conditions were $T_{\rm{surf}}$ = 165 K at a pressure of $10^{6}$ Ba (1 bar). Additional calculations for different surface temperatures are in Appendix B.
@@ 3.1 Differences due to the details of the adiabatic gradient
Although it may be tempting to use these results to select the ”best” method, in Appendix B we show that it depends on the surface temperature which method comes closest to the isentropic profile. These results demonstrate that the method of calculating the adiabatic gradient significantly influences the characterisation of giant planets interiors. We next examine how these results change when using the non-logarithmic form of the temperature equation.
@@ 4 Summary & Conclusions
Acknowledgements.
We thank Christopher Mankovich for helpful discussions. We acknowledge support from the Swiss National Science Foundation (SNSF) grant 200020_215634 and the National Centre for Competence in Research ‘PlanetS’ supported by SNSF. Extensive use was also made of the Python packages Jupyter (Kluyver et al. 2016), Matplotlib (Hunter 2007), NumPy (Harris et al. 2020), SciPy (Virtanen et al. 2020), planetsynth (Müller & Helled 2021) and tinyeos (Müller 2025).
@@ Appendix A The influence of the numerical solver on adiabatic interiors
In this section, we investigate how the choice of numerical solver for Equation 3 affects the adiabatic interiors in static interior models. As in Section 3, the differential equation was solved with scipy.integrate.solve_ivp, but we varied the integration methods. The adiabatic gradient was calculated from the Chabrier & Debras (2021) tables assuming an ideal mixture $S=XS_{X}+YS_{Y}+S_{\rm{mix}}$ and using Equation 5, where the derivatives were calculated from cubic splines.

### 20 Apr 2026 v2

# Resizing the giants: How modelling adiabatic interiors impacts predicted planetary radii

@@ Abstract
The interiors of giant planets are commonly assumed to be convective and adiabatic, making the adiabatic temperature gradient a key ingredient in interior and evolution models. Multiple numerically distinct methods exist for computing this gradient, yet their impact on inferred planetary structure and radius has not been systematically assessed. We investigate how the numerical treatment of adiabatic temperature profiles affects inferred planetary radii and internal structure, comparing different methods for evaluating the adiabatic gradient against a ground-truth isentropic baseline, for both the logarithmic and non-logarithmic forms of the temperature differential equation. Static interior models of a one Jupiter mass planet were computed using a state-of-the-art hydrogen-helium equation of state. The choice of numerical method significantly impacts the inferred interior structure and radius. Using the logarithmic temperature equation, central temperatures deviate by several thousand Kelvin and surface radii differ by up to 3.4%, exceeding the one-percent precision of current giant exoplanet radius measurements threefold. The non-logarithmic form reduces deviations to below $\sim$1% for most methods. We recommend spline derivatives to evaluate the adiabatic gradient, combined with the non-logarithmic temperature equation. Finite differencing and direct use of tabulated gradients or derivatives should be avoided.
keywords:
planets and satellites: general, gaseous planets, interiors – method: numerical
pubyear: 2026pagerange: Resizing the giants: How modelling adiabatic interiors impacts predicted planetary radii5
@@ 1 Introduction
Giant planets are key objects for understanding planetary system formation and evolution, and their interior structure provides crucial constraints on their composition, thermal state, and history (Guillot, 1999; Burrows et al., 2001; Fortney and Nettelmann, 2010; Thorngren and Fortney, 2019; Müller and Helled, 2023; Helled and Howard, 2026). Accurately characterising giant planet interiors therefore requires reliable structural models. A common and well-established assumption is that giant planet interiors are largely convective, leading to adiabatic (isentropic) temperature profiles (e.g., Stevenson, 1982; Guillot, 2005; Helled et al., 2014). The adiabatic temperature gradient thus plays a central role in both static interior models and long-term evolution calculations. However, despite being mathematically equivalent in the continuous limit, different numerical implementations of the adiabatic gradient can yield meaningfully different results, and the sensitivity of interior models to these numerical choices has not been systematically explored. In this work, we investigate and quantify the differences in the planetary interiors and radii that arise when using different methods of modelling adiabatic interiors of giant planets.
@@ 2 Methods
Giant planets are commonly modelled as a spherically symmetric objects in hydrostatic equilibrium. Under these assumptions the structure equations are (e.g.., Kippenhahn et al., 2012):
@@
where $r$, $m$, $\rho$, $p$ and $T$ are the radius, mass, density, pressure and temperature, G is the gravitational constant, and $\nabla_{\rm{rad,ad}}$ are the radiative and adiabatic temperature gradients. These equations are solved together with an appropriate equation of state and boundary conditions. Equation 3 further assumes the simple case [^1] While more sophisticated models use the mixing length theory (Böhm-Vitense, 1958) to determine the temperature gradient in convective regions, the adiabatic gradient still plays a key role (e.g., Kippenhahn et al., 2012). that the temperature gradient inside the planet is the minimum of the radiative or the adiabatic gradient, which is equivalent to the Schwarzschild criterion (Schwarzschild, 1906). In this work, we focus solely on fully convective interiors, such that the temperature gradient is always the adiabatic one. The form of $\nabla_{\rm{ad}}$ commonly used in stellar and planetary structure calculations is the logarithmic temperature derivative with respect to pressure at constant entropy $S$ (e.g., Lamers and Levesque, 2017):
@@
where $S$ is the entropy. Since the adiabatic gradient is a material property, calculating it requires an appropriate equation of state. The equation of state is generally given in a tabulated form using either the thermodynamic basis $(p,T)$ or $(\rho,T)$, and is commonly evaluated with cubic splines since derivatives can be required (for example, when using the Henyey method to solve the evolution equations (Henyey et al., 1965)). For simplicity and without loss of generality, we here focus on hydrogen-helium mixtures and do not include heavy elements.
The state-of-the-art hydrogen-helium equation of state for modelling giant planets and brown dwarfs are the tables from Chabrier and Debras (2021) (hereafter CD21). These tables provide the entropy, logarithmic entropy derivatives with respect to pressure and temperature, and the adiabatic gradient as a function of the pressure-temperature or density-temperature thermodynamic bases. There are tables for pure hydrogen and helium, or hydrogen-helium mixtures for a few values of the helium mass fraction. The hydrogen-helium mixture tables offer (among others [^2] Using Maxwell relations and the triple-product-rule, there are many ways deriving expressions for the adiabatic temperature gradient; see, e.g., Hansen and Kawaler (1994) or Equation 12 in Chabrier et al. (2019). These can be useful if the thermodynamic basis is not $(p,T)$. two relatively simple ways modelling adiabatic interiors: using the tabulated values for $\nabla_{\rm{ad}}$, or calculating it using the triple-product rule
@@ 3 Results
In this section we performed several numerical experiments with Python to determine the influence the previously mentioned numerical details have on the interiors of giant planets. We assumed a proto-solar hydrogen-helium mixture and evaluated the equations of state tables using cubic splines with scipy.interpolate.RegularGridInterpolator. The temperature equation was solved with scipy.integrate.solve_ivp using the explicit fifth-order Runge-Kutta method RK45 (additional solvers were tested in Appendix A), and the structure equations with a relaxation method (Thorngren et al., 2016; Mankovich et al., 2019; Mankovich and Fortney, 2020). We assumed a planetary mass of 1 MJ and the surface boundary conditions were $T_{\rm{surf}}$ = 165 K at a pressure of $10^{6}$ Ba (1 bar). Additional calculations for different surface temperatures are in Appendix B.
@@ 3.1 Differences due to the details of the adiabatic gradient
Although it may be tempting to use these results to select the "best" method, in Appendix B we show that it depends on the surface temperature which method comes closest to the isentropic profile. These results demonstrate that the method of calculating the adiabatic gradient significantly influences the characterisation of giant planets interiors. We next examine how these results change when using the non-logarithmic form of the temperature equation.
@@ 3.3 Practical recommendations
In this work we focused on the effect of the numerical details on the most widely used hydrogen-helium equation of state. We note that Howard et al. (2023) previously compared interpolation schemes and found that they affect the models, and also suggested that the pressure-temperature grid of the CD21 tables likely should be finer in particular regions of the parameter space. Furthermore, the details of the comparisons likely would have changed when other tables (e.g., Saumon et al., 1995; Becker et al., 2014). However, the main goal of the models presented here is to highlight that there are significant effects depending on the employed numerical scheme. Furthermore, our main recommendation to enforce an isentropic profile whenever possible should apply to all tabulated equations of state.
@@ 4 Summary & Conclusions
Acknowledgements
We thank Christopher Mankovich for helpful discussions. We acknowledge support from the Swiss National Science Foundation (SNSF) grant 200020_215634 and the National Centre for Competence in Research ‘PlanetS’ supported by SNSF. Extensive use was also made of the Python packages Jupyter (Kluyver et al., 2016), Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), planetsynth (Müller and Helled, 2021) and tinyeos (Müller, 2025).
Data Availability
The code and all interiors models in this work will be made available upon reasonable request to the authors.
@@ Appendix A The influence of the numerical solver on adiabatic interiors
In this section, we investigate how the choice of numerical solver for Equation 3 affects the adiabatic interiors in static interior models. As in Section 3, the differential equation was solved with scipy.integrate.solve_ivp, but we varied the integration methods. The adiabatic gradient was calculated from the Chabrier and Debras (2021) tables assuming an ideal mixture $S=XS_{X}+YS_{Y}+S_{\rm{mix}}$ and using Equation 5, where the derivatives were calculated from cubic splines.